# Solving PDEs in One Shot via Fourier Features with Exact Analytical Derivatives

### v1

@@ Abstract
Recent random feature methods for solving partial differential equations (PDEs) reduce computational cost compared to physics-informed neural networks (PINNs) but still rely on iterative optimization or expensive derivative computation. We observe that sinusoidal random Fourier features possess a cyclic derivative structure: the derivative of any order of $\sin(\mathbf{W}\cdot\mathbf{x}+b)$ is a single sinusoid with a monomial prefactor, computable in $O(1)$ operations. Alternative activations such as $\tanh$, used in prior one-shot methods like PIELM, lack this property: their higher-order derivatives grow as $O(2^{n})$ terms, requiring automatic differentiation for operator assembly. We propose FastLSQ, which combines frozen random Fourier features with analytical operator assembly to solve linear PDEs via a single least-squares call, and extend it to nonlinear PDEs via Newton–Raphson iteration where each linearized step is a FastLSQ solve. On a benchmark of 17 PDEs spanning 1 to 6 dimensions, FastLSQ achieves relative $L^{2}$ errors of $10^{-7}$ in 0.07 s on linear problems, three orders of magnitude more accurate and significantly faster than state-of-the-art iterative PINN solvers, and $10^{-8}$ to $10^{-9}$ on nonlinear problems via Newton iteration in under 9s.
@@ 1 Introduction
Random feature methods (RFMs) offer a middle ground. By approximating the solution as a linear combination of randomly sampled basis functions with frozen parameters, RFMs reduce the problem to fitting a linear output layer. For linear PDEs, this yields a linear system in the output coefficients for any choice of basis, as noted by PIELM (Dwivedi and Srinivasan, 2020) (using $\tanh$) and RF-PDE (Liao, 2024). However, RF-PDE still requires between 600 and 2000 epochs of iterative optimization, and even one-shot methods like PIELM must compute derivatives of the basis functions at every collocation point via automatic differentiation, a step whose cost depends critically on the activation choice.
@@
While any frozen-feature model yields a linear system for linear PDEs, the practical bottleneck is assembling the PDE operator matrix $A_{ij}=\mathcal{L}[\phi_{j}](\mathbf{x}_{i})$. For sinusoidal features $\phi_{j}(\mathbf{x})=\sin(\mathbf{W}_{j}\cdot\mathbf{x}+b_{j})$, we show that $\mathcal{L}[\phi_{j}]$ admits an exact closed-form expression for any linear differential operator of any order, requiring only $\mathcal{O}(1)$ operations per entry and no automatic differentiation. This is a consequence of the cyclic derivative structure of sinusoids: the $n$-th derivative cycles through $\{\sin,\cos,-\sin,-\cos\}$ with a monomial weight prefactor. Alternative bases such as $\tanh$ lack any comparable closed-form pattern (Proposition 2.2), so methods like PIELM must rely on automatic differentiation to assemble the operator matrix, incurring graph construction overhead and precluding the direct, graph-free evaluation that sinusoids enable.
@@
Figure 1: Architecture comparison. PINNs and RF-PDE require iterative optimization (top row). PIELM and Fast-LSQ both solve a linear system in one shot, but PIELM must invoke automatic differentiation to assemble the operator matrix from $\tanh$ features (middle row), whereas Fast-LSQ computes each operator entry analytically in $\mathcal{O}(1)$ (bottom row).
Figure 1 illustrates the key architectural differences. Both PIELM and Fast-LSQ are one-shot solvers that reduce to a single least-squares call, so neither involves iterative training. The distinction lies entirely in how the operator matrix $\mathbf{A}$ is assembled: PIELM evaluates $\mathcal{L}[\tanh(\mathbf{W}_{j}^{\top}\mathbf{x}+b_{j})]$ via automatic differentiation, while Fast-LSQ evaluates $\mathcal{L}[\sin(\mathbf{W}_{j}^{\top}\mathbf{x}+b_{j})]$ via a closed-form formula (§2.2). The practical consequences are: (1) no autodiff overhead—the operator matrix is assembled without constructing or traversing any computational graph; (2) superior spectral accuracy, with $10\times$ to $1000\times$ lower errors at equal feature counts (§3); and (3) deterministic output with no dependence on learning rate, optimizer state, or training schedule.
We make three contributions: First, we observe that the elementary cyclic derivative structure of sinusoids (2) enables graph-free, closed-form assembly of the PDE operator matrix for arbitrary linear differential operators, and we build on this to construct Fast-LSQ: a one-shot solver that assembles the operator matrix analytically and obtains the solution via a single least-squares call (§2.3). Second, we extend Fast-LSQ to nonlinear PDEs via Newton–Raphson iteration, where each linearized step reuses the analytical assembly, achieving $10^{-8}$ to $10^{-9}$ accuracy in under 9 s (§2.4). Third, on 17 PDEs (5 linear, 5 nonlinear solver-mode, 7 nonlinear regression) in up to 6 dimensions, Fast-LSQ achieves $10^{-7}$ in under 0.1 s on linear problems and $10^{-8}$ on nonlinear problems, outperforming all baselines by large margins (§3).
@@ 2.2 Exact Analytical Derivatives of Sinusoidal Features
The central structural property exploited by Fast-LSQ is the following.
The multivariate form follows immediately from the chain rule and the cyclic derivatives of $\sin$. For any multi-index $\alpha=(\alpha_{1},\ldots,\alpha_{d})\in\mathbb{N}_{0}^{d}$ with $|\alpha|=\sum_{k}\alpha_{k}$:
@@
where $\Phi_{0}=\sin$, $\Phi_{1}=\cos$, $\Phi_{2}=-\sin$, $\Phi_{3}=-\cos$. The identity itself is elementary; what matters is its consequence for PDE solving: for any linear differential operator $\mathcal{L}$ of any order, every entry of the operator matrix $A_{ij}=\mathcal{L}[\phi_{j}](\mathbf{x}_{i})$ reduces to a single trigonometric evaluation multiplied by a monomial in the weights $\mathbf{W}_{j}$, computable without automatic differentiation or computational graph construction.
@@ 2.3 Solver Mode: Direct Linear Solve
This linearity in $\boldsymbol{\beta}$ holds for any frozen-feature basis, including PIELM’s $\tanh$. The advantage of sinusoids is twofold: each entry $\mathcal{L}[\phi_{j}](\mathbf{x}_{i})$ is computed analytically without autodiff, and sinusoidal features yield superior spectral accuracy for smooth solutions (demonstrated empirically in §3). Algorithm 1 summarizes the complete procedure.
@@ 3.1 Setup
All methods use N=1500N{=}1500 features (3 blocks $\times$ 500), $M_{1}{=}10{,}000$ interior and $M_{2}{=}2{,}000$ boundary collocation points for linear PDEs ($M_{1}{=}5{,}000$ for Newton solver and regression modes). The boundary penalty is $\lambda{=}100$ except for Wave 2D-MS ($\lambda{=}1000$). Bandwidth $\sigma$ is selected via grid search over 9 values in $\{0.5,1,2,3,5,8,10,12,15\}$ with 3 trials. The Newton solver uses $\mu{=}10^{-10}$ Tikhonov regularization, backtracking line search, and up to 30 iterations (48 with continuation for Burgers). All experiments use PyTorch on a single NVIDIA GPU; the dominant cost is the $\mathcal{O}(MN^{2})$ least-squares solve. Reported wall-clock times include feature construction, matrix assembly, and the solve itself. All errors are evaluated out-of-sample on a separate dense test set of 5000 points drawn independently from the collocation points used during solving (different random seed). We report both the relative $L^{2}$ value error $\|u_{N}-u\|_{L^{2}}/\|u\|_{L^{2}}$ and the PDE residual norm $\|\mathcal{L}[u_{N}]-f\|_{L^{2}}$ evaluated on this independent test set; the latter confirms that low function error is not an artifact of overfitting the collocation points.
@@ 3.2 Linear PDE Results
Table 1: Results on linear PDEs (solver mode). Relative $L^{2}$ errors and wall-clock times. PINNacle times are the fastest across all 11 PINN variants from Table 12 of Zhongkai et al. (2024); PINNacle errors are the best across variants. “N/A” indicates the problem was not benchmarked in that work.
Fast-LSQ (sin) PIELM (tanh) PINNacle RF-PDE
Problem Time $L^{2}$ Err Time $L^{2}$ Err Time $L^{2}$ Err Time $L^{2}$ Err
Poisson 5D 0.07 s 4.8e-7 0.07 s 4.7e-6 $\sim$1780 s 4.7e-4 N/A 7.4e-4
@@
Table 1 shows the results for the five linear PDEs. On Poisson 5D, Fast-LSQ achieves a relative $L^{2}$ error of $4.8\times 10^{-7}$ in 0.07 s. This is three orders of magnitude more accurate than PINNacle’s best variant ($4.7\times 10^{-4}$) and approximately $1000\times$ better than RF-PDE ($7.4\times 10^{-4}$), while being roughly $25{,}000\times$ faster than the fastest PINNacle variant ($\sim$1780 s for the high-dimensional Poisson benchmark). On Wave 1D, where PINNs notoriously struggle ($9.8\times 10^{-2}$ best error), Fast-LSQ achieves $1.3\times 10^{-6}$, a five-order-of-magnitude accuracy improvement, in 0.06 s versus $\sim$272 s for PINNacle, a $4{,}500\times$ speedup. High-frequency Helmholtz (k=10k{=}10) and Maxwell validate spectral accuracy at $10^{-6}$$10^{-7}$ on problems for which no PINNacle benchmarks exist.
@@ 3.3 Nonlinear PDE Results: Newton Solver Mode
Table 2: Nonlinear PDEs solved via Newton–Fast-LSQ in true solver mode (no knowledge of the exact solution during solving). The regression column fits the known exact solution directly for comparison. All Newton runs use $\mu=10^{-10}$ Tikhonov regularization, backtracking line search, and warm-start from the linear part.
Newton–Fast-LSQ (Solver) Regression
Problem $\sigma$ Iters Time $L^{2}$ Err $|\nabla|$ Err Regr. $L^{2}$
NL-Poisson 2D ($u^{3}$) 3.0 30 8.2 s 6.1e-8 5.4e-7 1.9e-7
Bratu 2D ($\lambda{=}1$) 2.0 30 7.0 s 1.5e-7 8.8e-7 N/A
Steady Burgers ($\nu{=}0.1$) 10.0 48 7.4 s 3.9e-9 3.6e-9 3.3e-8
NL-Helmholtz 2D (k=3k{=}3) 5.0 30 8.8 s 2.4e-9 2.6e-8 N/A
Allen–Cahn 1D ($\varepsilon{=}0.1$) 3.0 30 4.2 s 6.0e-8 5.4e-7 1.2e-8
Table 2 demonstrates that Newton–Fast-LSQ extends the method to nonlinear PDEs with remarkable accuracy. The solver achieves between $10^{-8}$ and $10^{-9}$ relative $L^{2}$ errors on all five problems in under 9 s per problem, without any knowledge of the exact solution during solving. On NL-Poisson and Steady Burgers, Newton–Fast-LSQ outperforms the regression baseline that fits the exact solution directly ($6.1\times 10^{-8}$ vs. $1.9\times 10^{-7}$ and $3.9\times 10^{-9}$ vs. $3.3\times 10^{-8}$, respectively). This counter-intuitive result occurs because the Newton solver enforces the PDE structure as an additional constraint beyond pure function fitting: the physics provides regularization that the regression mode lacks. On Allen-Cahn, the regression baseline achieves a slightly lower error ($1.2\times 10^{-8}$ vs. $6.0\times 10^{-8}$), suggesting that the Newton iteration has not fully converged for this stiff problem. We emphasize that all reported errors are computed on an independent test set, not on the collocation points used during solving, and that the PDE residual norms on these test points (Table 2, $|\nabla|$ column) corroborate the value errors.
Table 3: Newton–Fast-LSQ runtime breakdown. Time per iteration is computed as total time divided by total Newton iterations (including continuation steps for Burgers). The dominant cost per iteration is the $\mathcal{O}(MN^{2})$ Tikhonov-regularized least-squares solve.
Problem $\sigma$ Newton iters Continuation steps Time/iter Total time
NL-Poisson 2D ($u^{3}$) 3.0 30 0.27 s 8.2 s
Bratu 2D ($\lambda{=}1$) 2.0 30 0.23 s 7.0 s
Steady Burgers ($\nu{=}0.1$) 10.0 48 4 0.15 s 7.4 s
NL-Helmholtz 2D (k=3k{=}3) 5.0 30 0.29 s 8.8 s
Allen–Cahn 1D ($\varepsilon{=}0.1$) 3.0 30 0.14 s 4.2 s
Table 3 provides a detailed runtime breakdown. Each Newton iteration costs between 0.14 s and 0.29 s, with the variation driven by problem dimensionality: the 2D problems (NL-Poisson, Bratu, NL-Helmholtz) are more expensive because the feature matrices are larger and the Jacobian assembly involves more collocation points. The 1D problems (Burgers, Allen–Cahn) are cheaper per iteration; Burgers requires more total iterations (48 across 4 continuation steps) but each step is fast, yielding a comparable total runtime. The continuation strategy is essential for the advection-dominated Burgers problem: it divides the 48 iterations across 4 stages ($\nu=1.0,0.5,0.2,0.1$), with roughly 12 iterations per stage, enabling convergence where direct Newton iteration at $\nu=0.1$ fails to converge.
The convergence profiles (Appendix B) reveal two distinct regimes: a rapid initial phase where the residual drops by several orders of magnitude (typically in 2–3 iterations), followed by a slow plateau phase where the line search is frequently active ($\alpha<1$) and the residual decreases by less than one order of magnitude over the remaining iterations. This plateau behavior is characteristic of the Tikhonov-regularized solve: the regularization parameter $\mu=10^{-10}$ prevents the Newton step from fully resolving the Jacobian near-nullspace, trading convergence rate for robustness.

### v3

# FastLSQ: Solving PDEs in One Shot via Fourier Features with Exact Analytical Derivatives

@@ Abstract
We present FastLSQ, a framework for PDE solving and inverse problems built on trigonometric random Fourier features with exact analytical derivatives. Trigonometric features admit closed-form derivatives of any order in $\mathcal{O}(1)$, enabling graph-free operator assembly without autodiff. Linear PDEs: one least-squares call; nonlinear: Newton–Raphson reusing analytical assembly. On 17 PDEs (1–6D), FastLSQ achieves $10^{-7}$ in 0.07 s (linear) and $10^{-8}$$10^{-9}$ in <<9 s (nonlinear), orders of magnitude faster and more accurate than iterative PINNs. Analytical higher-order derivatives yield a differentiable digital twin; we demonstrate inverse problems (heat-source, coil recovery) and PDE discovery. Code: github.com/sulcantonin/FastLSQ; pip install fastlsq.
@@ 1 Introduction
Random feature methods (RFMs) offer a middle ground. By approximating the solution as a linear combination of randomly sampled basis functions with frozen parameters, RFMs reduce the problem to fitting a linear output layer. For linear PDEs, this yields a linear system in the output coefficients for any choice of basis, as noted by PIELM (Dwivedi and Srinivasan, 2020) (using $\tanh$) and RF-PDE (Liao, 2024). However, RF-PDE still requires between 600 and 2000 epochs of iterative optimization. One-shot methods like PIELM use $\tanh$ activations, which lack a closed-form cyclic derivative structure; PIELM therefore relies on manual, problem-specific symbolic calculus to assemble $\mathcal{L}[\tanh]$ for each PDE operator.
@@
While any frozen-feature model yields a linear system for linear PDEs, the practical bottleneck is assembling the PDE operator matrix $A_{ij}=\mathcal{L}[\phi_{j}](\mathbf{x}_{i})$. For sinusoidal features $\phi_{j}(\mathbf{x})=\sin(\mathbf{W}_{j}\cdot\mathbf{x}+b_{j})$, we show that $\mathcal{L}[\phi_{j}]$ admits an exact closed-form expression for any linear differential operator of any order, requiring only $\mathcal{O}(1)$ operations per entry and no automatic differentiation. This is a consequence of the cyclic derivative structure of sinusoids: the $n$-th derivative cycles through $\{\sin,\cos,-\sin,-\cos\}$ with a monomial weight prefactor. Alternative bases such as $\tanh$ lack any comparable closed-form pattern (Proposition 2.2), so methods like PIELM require manual, problem-specific derivation of $\mathcal{L}[\tanh]$ for each new PDE operator; there is no universal formula. Fast-LSQ’s operator-agnostic advantage is that one formula (2) applies to any $\mathcal{L}$.
@@
Figure 1: Architecture comparison. PINNs and RF-PDE require iterative optimization (top row). PIELM and Fast-LSQ both solve a linear system in one shot, but PIELM requires handwritten, problem-specific calculus to assemble $\mathcal{L}[\tanh]$ for each PDE operator (middle row), whereas Fast-LSQ uses a single operator-agnostic formula for any $\mathcal{L}$ (bottom row).
[…]
We make three contributions: First, we observe that the elementary cyclic derivative structure of sinusoids (2) enables graph-free, closed-form assembly of the PDE operator matrix for arbitrary linear differential operators, and we build on this to construct Fast-LSQ: a one-shot solver that assembles the operator matrix analytically and obtains the solution via a single least-squares call (§2.3). Second, we extend Fast-LSQ to nonlinear PDEs via Newton–Raphson iteration, where each linearized step reuses the analytical assembly, achieving $10^{-8}$ to $10^{-9}$ accuracy in under 9 s (§2.4). Third, on 17 PDEs (5 linear, 5 nonlinear solver-mode, 7 nonlinear regression) in up to 6 dimensions, Fast-LSQ achieves $10^{-7}$ in under 0.1 s on linear problems and $10^{-8}$ on nonlinear problems, outperforming all baselines by large margins (§3). Furthermore, we run a comparison with RBF Kansa, scikit-fem P2 FEM, and scipy.solve_bvp on all 10 PDEs, demonstrating that Fast-LSQ is the only method applicable to every problem, including high-dimensional ($d{\geq}5$) and hyperbolic space-time equations for which conventional solvers have no applicable path.
@@ 2.2 Exact Analytical Derivatives of Sinusoidal Features
The central structural property exploited by Fast-LSQ is the following. The multivariate form follows immediately from the chain rule and the cyclic derivatives of $\sin$. For any multi-index $\alpha=(\alpha_{1},\ldots,\alpha_{d})\in\mathbb{N}_{0}^{d}$ with $|\alpha|=\sum_{k}\alpha_{k}$:
@@
where $\Phi_{0}=\sin$, $\Phi_{1}=\cos$, $\Phi_{2}=-\sin$, $\Phi_{3}=-\cos$. The identity itself is elementary; what matters is its consequence for PDE solving: for any linear differential operator $\mathcal{L}$ of any order, every entry of the operator matrix $A_{ij}=\mathcal{L}[\phi_{j}](\mathbf{x}_{i})$ reduces to a single trigonometric evaluation multiplied by a monomial in the weights $\mathbf{W}_{j}$, computable without automatic differentiation or computational graph construction. This analytical formulation also enables a differentiable trick: the bandwidth $\sigma$ (or an anisotropic covariance) can be optimised via gradient descent, since gradients flow through the exact inner solve (Appendix E).
@@ 2.3 Solver Mode: Direct Linear Solve
This linearity in $\boldsymbol{\beta}$ holds for any frozen-feature basis, including PIELM’s $\tanh$. The advantages of sinusoids are: (1) operator-agnostic assembly—one formula (2) for any $\mathcal{L}$, versus PIELM’s manual derivation per operator; (2) superior spectral accuracy for smooth and oscillatory solutions (§3); and (3) high-fidelity gradients, because derivatives are phase-shifted sinusoids scaled by $\mathcal{O}(\sigma)$, so gradient errors stay within one order of magnitude of value errors (§3.6). Algorithm 1 summarizes the complete procedure.
@@ 2.4 Newton–Raphson Extension for Nonlinear PDEs
2.5 Differentiable System and Hyperparameter Optimisation
The solution and its derivatives are fully differentiable with respect to design parameters, boundary conditions, and source terms. Because the forward pass is a single least-squares solve (or a short Newton sequence), the entire system is cheap to optimise: gradients flow through the pre-factored linear solve. Just as we use Newton–Raphson to optimise the coefficients $\boldsymbol{\beta}$ for nonlinear PDEs, we can use L-BFGS-B to optimise hyperparameters of unknown equations. We demonstrate this in §3.8: inverse heat-source localisation (recovering source positions from sparse sensor data), sparse-sensor coil localisation in electrostatics, and PDE synthesis from data via analytical derivative dictionaries.
@@ 3.1 Setup
All methods use N=1500N{=}1500 features (3 blocks $\times$ 500), $M_{1}{=}10{,}000$ interior and $M_{2}{=}2{,}000$ boundary collocation points for linear PDEs ($M_{1}{=}5{,}000$ for Newton solver and regression modes). The boundary penalty is $\lambda{=}100$ except for Wave 2D-MS ($\lambda{=}1000$). Bandwidth $\sigma$ is selected via grid search over 9 values in $\{0.5,1,2,3,5,8,10,12,15\}$ with 3 trials. The Newton solver uses $\mu{=}10^{-10}$ Tikhonov regularization, backtracking line search, and up to 30 iterations (48 with continuation for Burgers). All experiments use PyTorch on a single NVIDIA T4 GPU (16 GB). Runtime is directly proportional to the number of collocation points $M$ and the square of the number of features $N$. Reported wall-clock times include feature construction, matrix assembly, and the solve itself. All errors are evaluated out-of-sample on a separate dense test set of 5000 points drawn independently from the collocation points used during solving (different random seed). We report both the relative $L^{2}$ value error $\|u_{N}-u\|_{L^{2}}/\|u\|_{L^{2}}$ and the PDE residual norm $\|\mathcal{L}[u_{N}]-f\|_{L^{2}}$ evaluated on this independent test set; the latter confirms that low function error is not an artifact of overfitting the collocation points.
@@ 3.2 Linear PDE Results
Table 1: Results on linear PDEs (solver mode). Relative $L^{2}$ errors and wall-clock times. PINNacle times are fastest across all 11 variants from Table 12 of Zhongkai et al. (2024). “N/A” = not benchmarked in that work; “—” = not applicable to this problem class. RF-PDE does not benchmark our exact linear problems; the closest available result is nonlinear Poisson in d=4d{=}4 (51 s, 1000 epochs) and d=8d{=}8 (38 s, 1500 epochs) from Table 3 of Liao (2024); we report the d=4d{=}4 timing as the best available lower bound for 5-D. Conventional baseline: scikit-fem P2 FEM at 16 641 DoF for Helmholtz 2D (the only 2-D elliptic spatial problem in this group). Poisson 5D and Heat 5D are 5-/6-dimensional; FEM mesh size grows as $h^{-d}$, making $d{\geq}5$ computationally intractable. Wave 1D and Maxwell 2D TM are posed as space-time problems with hyperbolic structure; scikit-fem targets elliptic spatial problems only, and scipy.solve_bvp handles steady 1-D BVPs only.
Fast-LSQ (sin) PIELM (tanh) PINNacle RF-PDE Conv.
Problem Time $L^{2}$ Time $L^{2}$ Time $L^{2}$ Time $L^{2}$ Time $L^{2}$
Poisson 5D 0.07 s 4.8e-7 0.07 s 4.7e-6 $\sim$1780 s 4.7e-4 $\sim$51 s 7.4e-4
@@
Table 1 shows the results for the five linear PDEs. On Poisson 5D, Fast-LSQ achieves a relative $L^{2}$ error of $4.8\times 10^{-7}$ in 0.07 s. This is three orders of magnitude more accurate than PINNacle’s best variant ($4.7\times 10^{-4}$) and approximately $1000\times$ better than RF-PDE ($7.4\times 10^{-4}$), while being roughly $25{,}000\times$ faster than the fastest PINNacle variant ($\sim$1780 s) and $700\times$ faster than RF-PDE ($\sim$51 s for 1000 epochs of iterative optimisation on the closest available benchmark). On Wave 1D, where PINNs notoriously struggle ($9.8\times 10^{-2}$ best error), Fast-LSQ achieves $1.3\times 10^{-6}$, a five-order-of-magnitude accuracy improvement, in 0.06 s versus $\sim$272 s for PINNacle, a $4{,}500\times$ speedup. High-frequency Helmholtz (k=10k{=}10) and Maxwell validate spectral accuracy at $10^{-6}$$10^{-7}$ on problems for which no PINNacle benchmarks exist. For Helmholtz 2D—the only 2-D elliptic problem in this group and the one where a conventional FEM solver applies—scikit-fem P2 achieves $4.0\times 10^{-5}$ using 16 641 DoF; Fast-LSQ surpasses this accuracy ($1.9\times 10^{-6}$) with only 1500 features. For the remaining four problems (Poisson 5D, Heat 5D, Wave 1D, Maxwell 2D TM) no grid-based conventional solver is applicable due to the high dimensionality or hyperbolic space-time structure.
@@ 3.3 Nonlinear PDE Results: Newton Solver Mode
Table 2: Nonlinear PDEs via Newton–Fast-LSQ (true solver mode, no access to exact solution). All runs use $\mu=10^{-10}$ Tikhonov, backtracking line search, and warm-start. Time/iter is the dominant $\mathcal{O}(MN^{2})$ cost. Burgers uses 4-stage continuation ($\nu=1.0\to 0.1$). Conventional baseline $L^{2}$ error: scikit-fem P2 FEM (${\approx}4000$ DoF) for 2-D problems; scipy.integrate.solve_bvp (adaptive collocation) for 1-D problems.
Newton–Fast-LSQ Regr. Conv.
Problem $\sigma$ Iters Time/iter $L^{2}$ $|\nabla|$ Total $L^{2}$ $L^{2}$
NL-Poisson ($u^{3}$) 3 30 0.27 s 6.1e-8 5.4e-7 8.2 s 1.9e-7 2.6e-6
Bratu ($\lambda{=}1$) 2 30 0.23 s 1.5e-7 8.8e-7 7.0 s N/A 2.6e-6
Burgers ($\nu{=}0.1$) 10 48 0.15 s 3.9e-9 3.6e-9 7.4 s 3.3e-8 1.8e-10
NL-Helm. (k=3k{=}3) 5 30 0.29 s 2.4e-9 2.6e-8 8.8 s N/A 2.4e-6
Allen–Cahn ($\varepsilon{=}0.1$) 3 30 0.14 s 6.0e-8 5.4e-7 4.2 s 1.2e-8 1.2e-10
[…]
The convergence profiles (Appendix C) reveal two distinct regimes: a rapid initial phase where the residual drops by several orders of magnitude (typically in 2–3 iterations), followed by a slow plateau phase where the line search is frequently active ($\alpha<1$) and the residual decreases by less than one order of magnitude over the remaining iterations. This plateau behavior is characteristic of the Tikhonov-regularized solve: the regularization parameter $\mu=10^{-10}$ prevents the Newton step from fully resolving the Jacobian near-nullspace, trading convergence rate for robustness.